Quantitative Problem: Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis. Both projects' after-tax cash flows are shown on the time line below. Depreciation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firm's average project. Bellinger's WACC is 7%.

0 1 2 3 4
Project A -950 600 450 250 300
Project B -950 200 385 400 750

1. What is Project A's payback? Round your answer to four decimal places. Do not round your intermediate calculations.

2. What is Project A's discounted payback? Round your answer to four decimal places. Do not round your intermediate calculations.

3. What is Project B's payback? Round your answer to four decimal places. Do not round your intermediate calculations.

4.What is Project B's discounted payback? Round your answer to four decimal places. Do not round your intermediate calculations.

Please Help!!Thank you.
Answer #1
1
Project A
Year Cash flow stream Cumulative cash flow
0 -950 -950
1 600 -350
2 450 100
3 250 350
4 300 650
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 1 and 2
therefore by interpolation payback period = 1 + (0-(-350))/(100-(-350))
1.7778 Years
2
Project A
Year Cash flow stream Cumulative cash flow
0 -950 -950
1 600 -350
2 450 100
3 250 350
4 300 650
Discounted payback period is the time by which discounted cashflow cover the intial investment outlay
this is happening between year 1 and 2
therefore by interpolation payback period = 1 + (0-(-389.25))/(3.8-(-389.25))
1.9903 Years
Where
Discounting factor =(1 + discount rate)^(corresponding year)
Discounted Cashflow=Cash flow stream/discounting factor
3
Project B
Year Cash flow stream Cumulative cash flow
0 -950 -950
1 200 -750
2 385 -365
3 400 35
4 750 785
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 2 and 3
therefore by interpolation payback period = 2 + (0-(-365))/(35-(-365))
2.9125 Years
4
Project B
Year Cash flow stream Cumulative cash flow
0 -950 -950
1 200 -750
2 385 -365
3 400 35
4 750 785
Discounted payback period is the time by which discounted cashflow cover the intial investment outlay
this is happening between year 3 and 4
therefore by interpolation payback period = 3 + (0-(-100.29))/(471.88-(-100.29))
3.1753 Years
Where
Discounting factor =(1 + discount rate)^(corresponding year)
Discounted Cashflow=Cash flow stream/discounting factor